### v1

# Unveiling Inefficiencies in LLM-Generated Code: Toward a Comprehensive Taxonomy

Abstract
Large Language Models (LLMs) are widely adopted for automated code generation with promising results. Although prior research has assessed LLM-generated code and identified various quality issues- such as redundancy, poor maintainability, and sub-optimal performance- a systematic understanding and categorization of these inefficiencies remain unexplored. Without such knowledge, practitioners struggle to optimize LLM-generated code for real-world applications, limiting its adoption. This study can also guide improving code LLMs, enhancing the quality and efficiency of code generation. Therefore, in this study, we empirically investigate inefficiencies in LLM-generated code by state-of-the-art models, i.e., CodeLlama, DeepSeek-Coder, and CodeGemma. To do so, we analyze 492 generated code snippets in the HumanEval++ dataset. We then construct a taxonomy of inefficiencies in LLM-generated code that includes 5 categories (General Logic, Performance, Readability, Maintainability, and Errors) and 19 subcategories of inefficiencies. We then validate the proposed taxonomy through an online survey with 58 LLM practitioners and researchers. Our study indicates that logic and performance-related inefficiencies are the most popular, relevant, and frequently co-occur and impact overall code quality inefficiency. Our taxonomy provides a structured basis for evaluating the quality LLM-generated code and guiding future research to improve code generation efficiency.
Index Terms:
LLMs, Code Generation, Code Quality
I Introduction
Large Language Models (LLMs) have revolutionized software development by enabling automated code generation [1, 2, 3]. In recent years, we have witnessed the emergence of LLMs specifically tailored for code generation, including open-source models such as CodeLlama [4] and proprietary models like GPT-4 [5]. These models have demonstrated the ability to generate code across multiple programming languages, including Python, Java, and C/C++ [6], showcasing their potential to assist developers in various tasks [7, 8].
While LLMs show promising results in code generation, their generated code, similar to human-written one, exhibit various code quality issues [9, 8], limiting their seamless integration into production environments. For instance, LLM-generated code have shown to be suffered from redundancies, unnecessary computations, and suboptimal implementations [10], resulting in increased execution time, higher memory consumption, maintainability challenges, and import errors [11, 12]. Previous studies have also shown that LLMs may generate incorrect logic, misinterpret task requirements, and struggle to handle corner cases, leading to reliability concerns [13]. Even when functionally correct, LLM-generated code can face quality issues [14], being difficult to maintain [11], or unnecessarily complex [10], prone to security vulnerability and requiring manual intervention before deployment. These issues pose significant risks to software quality, particularly for developers adopting LLMs in real-world settings [15].
While prior research has examined specific quality aspects such as correctness, security, and maintainability in LLM-generated code [16, 17, 14], inefficiencies as a broader concept remain largely unexplored. Existing studies focus only on buggy-generated code [13, 14] to categorize bugs and identify root causes. However, this narrow focus overlooks a critical issue: even functionally correct code can exhibit inefficiencies that degrade software quality, such as performance bottlenecks [10], limiting real-world adoption of LLM-generated code.
In this study, we aim to bridge this gap by identifying, categorizing, and analyzing overall inefficiency patterns, regardless of correctness, in LLM-generated code from three leading open-source LLMs for code generation. We define inefficiencies as any issues in LLM-generated code that degrade its quality, hinder integration into larger systems, or even affect functionality in isolation. Understanding inefficiencies guides the development of more efficient LLMs and helps practitioners and researchers to improve LLM-generated code. Our investigation is done under the following Research Questions (RQs):
  • RQ1 What inefficiency patterns occur in LLM-generated code?
  • RQ2 How relevant are these inefficiencies to software practitioners and researchers?
We conducted our study using HumanEval+  a widely used benchmark for code generation composed of 164 human-crafted Python tasks [6]. We selected three leading open-source LLM models -CodeLlama, DeepSeek-Coder, and CodeGemma- to generate code for the different tasks, resulting in 492 generated code snippets. To assess LLM-generated code quality and requirements conformity, we employed a higher-capacity GPT-4o-mini model as a judge. The judgments provided initial guidelines for our manual exploratory and qualitative analysis to identify inefficiency patterns in LLM-generated code. Our study results in a taxonomy spanning 5 categories General Logic, Performance, Readability, Maintainability, and Errors, comprising 19 subcategories that is validated by surveying 58 software practitioners and researchers using LLM for coding assistance.
Our findings indicate that General Logic and Performance related inefficiencies are the most frequent, relevant and often co-occur with Errors, Maintainability and Readability and impact the overall efficiency of the code, suggesting that improving LLM reasoning and optimization capabilities could enhance overall code quality. Our survey confirmed these findings, with participants prioritizing correctness, but also expressing concerns about inefficiencies that increase debugging and refactoring efforts.
For LLM developers, our taxonomy offers insights into existing inefficiencies which can inform future targeted improvement directions. For researchers, our categorization provides a structured foundation for further studies on inefficiencies in LLM-generated code. Additionally, our findings can support the development of automated inefficiency detection and refactoring tools, which can enhance code quality as LLM-generated code becomes increasingly adopted in production systems. To summarize, this paper makes the following contributions:
  • We study inefficiencies observed in LLM-generated code and their prevalence.
  • We propose a taxonomy of inefficiencies in LLM-generated code based on their characteristics.
  • We validate our findings using an online survey and provide insights for researchers and practitioners.
  • We make our data and results publicly available in our replication package. [^1] https://anonymous.4open.science/r/LLMs_Code_Inefficiencies-EEB8
The rest of the paper is organized as follows: In Section II, we present the background. Next, we detail the methodology used to construct and validate the taxonomy of inefficiencies in Section III. In Section IV, we report our empirical findings presenting our taxonomy. Then, we discuss related work and threats in Section V and VI, respectively. Finally, Section VII concludes the paper.
II Background
In this section, we introduce the used three LLMs for code generation in our study, and describe the open coding methodology followed to construct the taxonomy of inefficiencies.
II-A LLMs for Code Generation.
Various LLMs have been introduced for automatic code generation; in our study, we focus on three models as described as follows. CodeLlama is an open-source family of models based on Llama 2 [18] from Meta, designed to support large input contexts and zero-shot instruction following, supporting coding tasks including code completion and infilling on surrounding content. The CodeLlama family includes the foundation model (CodeLlama), a Python-specialized variant (CodeLlama-Python), and an instruction-following model (CodeLlama-Instruct). Each of these models is available in different sizes (7B, 13B, 34B, and 70B parameters). This LLM is primarily trained on a near-deduplicated dataset of publicly available code, with 8% sourced from natural language datasets related to code, including discussions, questions, and answers, while a small portion is sampled to retain natural language understanding skills [4]. In our study, we focused on code completion and utilized the 7B parameter CodeLlama foundation model.
DeepSeek-Coder owns a range of open-source models trained from scratch on 2 trillion tokens across 87 programming languages. The pretraining data is organized at the repository level to enhance model capability. The models employ two key techniques: fill-in-the-middle and next-token prediction. Just like CodeLlama, DeepSeek-Coder is also available in different sizes, ranging from 1.3B to 33B parameters, with both base and instruction-tuned versions. These models leverage a 16K context window to improve code generation and infilling [19]. In our study, we focused on code completion and utilized the 6.7B parameter DeepSeek-Coder base model.
CodeGemma is a collection of specialized open-code models built on top of Google DeepMind’s Gemma models [20], further trained on 500 to 1000 billion tokens, primarily for code-related tasks such as code completion and generation. The CodeGemma family includes three models: a 7B pre-trained model, a 7B instruction-tuned model, and a specialized 2B model designed for code infilling and open-ended generation. The 2B model is trained entirely on code, while the 7B models are trained on a mix of 80% code and 20% natural language, sourced from deduplicated publicly available code repositories. These models are trained using a fill-in-the-middle task to enhance their coding capabilities [21]. In our study, we used the CodeGemma 7B pre-trained model for code completion.
II-B Open Coding.
Open coding is a qualitative research method used to analyze textual data, for example, survey responses and codes, by breaking it down into concepts and categories in an unstructured approach [22]. It consists of three parts: noticing, collecting, and thinking [23]. Initially, researchers examine closely raw data by looking into similarities and dissimilarities, notice patterns, and mark them with appropriate code, i.e. label to define categories based on noticed characteristics and properties. This approach is not linear as researchers often need to revisit data and adjust labels -adding, removing, or merging- to develop a final codebook. Open coding is known to be postformed code since codes are formed incrementally during the process [22]. Postformed codes have proven to be more useful in SE research than preformed labels [24, 25]. During open coding, independent reviewers begin by coding the same set of samples, identifying patterns and categories relevant to the study’s focus. They then engage in discussions to consolidate their findings into an initial codebook, resolving disagreements through communication or by involving a third party. To ensure reliability, the level of agreement between reviewers is measured before discussion [22]. Previous studies in software engineering have employed open coding for qualitative research, including the construction of taxonomies as we do here [26, 24, 25].
III Methodology
Figure 1: The methodology we followed for our study.
This section outlines the methodology we used in this paper to construct and validate the taxonomy of inefficiencies in LLM-generated code, consisting of 5 steps, as illustrated in Figure 1. First, we collected data by generating code using the HumanEval+ benchmark [6] with three open-source LLM4Code models: CodeLlama [4], DeepSeek-Coder [19], and GodeGemma [21]. Second, we leveraged a higher-capacity model (GPT-4o-mini) to judge the generated code and identify existing inefficiencies. In the third and fourth steps, we manually analyzed the code samples and the model’s judgments to develop a comprehensive taxonomy. Finally, in the fifth step, we surveyed practitioners to validate the LLM-generated code inefficiencies in our taxonomy. We describe details of each step in the rest of this section.
III-A Step 1: Data Selection and Collection
Regarding the data collection, considering the popularity and wide adoption of Python as a programming language[27], this study focused on Python-generated LLM code. Given the focus on Python, we selected HumanEval+ [6] (denoted as $\mathit{A}$ in Figure 1), a widely recognized and extensively employed benchmark for code generation [28, 29, 30, 31]. HumanEval+ enhances the original HumanEval [32]dataset by augmenting it with additional test cases, enabling a more rigorous and reliable assessment of functional correctness [6]. Such a benchmark comprises 146 Python programming tasks, each accompanied by (i) a task description provided as a doc string and (ii) associated unit tests to rigorously evaluate the correctness of the code snippets. Regarding the LLMs, we aimed to select those trained or fine-tuned for code generation tasks to enable a more meaningful and targeted analysis of inefficiencies. To ensure accessibility and feasibility, we established two selection criteria: (i) the models must be open source and (ii) have fewer than 8B parameters. Based on these criteria, we selected the three highest-performing models according to the HumanEval+ leaderboard [33] [^2] Accessed July 10, 2024: CodeLlama-7B [^3] https://huggingface.co/meta-llama/CodeLlama-7b-hf, GodeGemma-7B [^4] https://huggingface.co/google/codegemma-7b, and DeepSeek Coder-6.7B [^5] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base.
Once defined the setup for our study, we moved forward by generating the code snippets. Regarding the hyperparameters, we adopted a conservative approach, setting the temperature to 00, prioritizing deterministic outputs, and reducing randomness in code generation. For prompt construction, we adhered to default configuration from official Hugging Face documentation. This approach ensured that the models could generate code efficiently and accurately within the defined constraints. As a result of this step, we have a dataset of $164$ generated code snippets for each LLM, resulting in $492$ code snippets ($\mathit{B}$), from the initial dataset of tasks ($\mathit{A}$).
III-B Step 2: Judging LLM-generated Code
LLMs have been effectively used to assess code or tests [34, 35], proposing a paradigm where LLMs serve as judges (LaaJ). Prior studies have also leveraged LLM to assist in identifying a generated-code bug’s root causes [14]. In this context, we propose leveraging LLMs to assess the quality of generated code. Specifically, we employed the GPT-4o-mini model, which offers strong performance comparable to GPT-4 while being more cost-effective [36]. Regarding the hyperparameters, we considered a temperature of $\mathit{0}$ aiming to achieve a more conservative generation and max_tokens set to $\mathit{5120}$ to ensure sufficient space for detailed output. Knowing that GPT-4 models can be shaped and optimized by setting system prompts, which define the model’s persona [37], we adopted a persona of an expert in code quality analysis, with a focus on identifying specific inefficiencies in machine-generated code. Such a persona guides the model prioritizes practical and significant code inefficiencies rather than minor or irrelevant issues. Finally, we defined a fixed user prompt used as input when asking for judgment support (judger). In this prompt, we input the generated code snippet and its task description, providing a context to aid the judger in identifying logical inefficiencies in the generated code. As a result of this step, we enriched the dataset $\mathit{B}$ with GPT-generated judgments, creating dataset $\mathit{C}$, as denoted in Figure 1. Specifically, $\mathit{C}$ consists of 492 generated code samples, each paired with its corresponding LLM-generated judgment.
III-C Step 3 and 4: Manual Labeling
Step 3: Manual Labeling - Pilot Study With no predefined categories in previous studies, our objective is to construct a taxonomy of inefficiencies using an inductive, bottom-up approach through manual analysis of LLM-code generated snippets. To initiate this process, we randomly selected 196 samples $\mathit{C}$ (40%, pilot set), ensuring an equal proportion of cases from each model. The first two authors, senior Master’s student and a Postdoctoral Fellow with experience in Python and SE research, collaboratively analyzed and labeled the pilot set (40% of samples) using open coding [22]. The analysis involved examining the task description, the generated code, code quality judgment, and the execution and testing status (from the unit tests available in HumanEval+). Without predefined assumptions, inefficiencies were iteratively identified and grouped into categories and subcategories. Knowing that LLMs as judgers may hallucinate [35], we used their judgments as a starting point but independently verified and expanded the inefficiencies. During this process, the labelers assigned brief, descriptive labels to each inefficiency observed in the pilot set. These initial labels were then categorized, forming a hierarchical taxonomy of inefficiency types. This iterative process involved continuously refining the categories as they moved back and forth between individual samples and emerging patterns, ensuring that the taxonomy accurately reflected the inefficiencies underlying generated code. All tasks in the pilot set were thoroughly discussed, and the labelers reached a consensus on both categories and subcategories, resulting in a taxonomy comprising $5$ inefficiency categories and $18$ subcategories and a fully labeled pilot set. This step resulted a codebook of inefficiency patterns, with multi labeling allowed as a single code snippet can exhibit multiple inefficiencies [13].
Step 4: Taxonomy Refinement & Full Dataset Labeling After labeling the pilot set and constructing the initial codebook of inefficiencies, the labelers independently labeled the remaining 60% samples, following the same procedure. If any generated code could not be categorized within the existing taxonomy, they were marked Pending and reviewed to determine if new categories or subcategories should be created to accommodate these cases. Discussions, led to adding Syntax Error as a subcategory in the Errors category, resulting in a final taxonomy of 5 categories and 19 subcategories to better capture inefficiencies observed in LLM-generated code. Once all samples were labeled, we assessed inter-rater agreement to evaluate the consistency between the two labelers. Following similar studies [26, 13], we used Cohen’s Kappa [38], calculated at category and subcategory levels. Inter-rater agreement (Cohen’s Kappa) reached 0.846 (category) and 0.735 (subcategory), aligning with prior SE studies[26, 13]. Posts labeled as ‘Pending‘ were included in the computation of Cohen’s Kappa only when both raters agreed on this designation. Labeling disagreements were resolved through discussion, leading to a consensus on the final labeling.
III-D Step 5: Participants feedback and survey analysis
To evaluate the frequency and relevance of the identified inefficiencies, we conducted a survey targeting software practitioners and researchers using LLMs for code generation. Following prior studies [39, 13], we collected email addresses of GitHub users who collaborated on repositories containing LLM-generated code. To identify relevant repositories, we searched GitHub for repositories where at least one file contained code generated by LLM-based tools or models, including Copilot, ChatGPT, Codex, CodeLlama, Llama, and Llama-2. Specifically, we used the following keyword-based queries: association by/with with Copilot, ChatGPT, Codex, CodeLlama, Llama, Llama-2. This process yielded $411$ repositories. Next, we extracted publicly available email addresses of contributors from these repositories using PyDriller [40]. This resulted in $835$ unique email addresses, of which $51$ were unreachable, leaving us with $784$ successfully delivered emails.  [^6] Date of email mining: November 18, 2024. First, the survey outlined a detailed message about its purpose, scope, and estimated completion time. It was divided into two main parts. The first part includes demographic and open-ended questions regarding LLMs participants had used and related inefficiencies they had encountered based on their experience. The second part evaluated the frequency and relevance of the inefficiencies in LLM-generated code, previously defined in Step 4. For each inefficiency category, we provided a general description, followed by their specific subcategories, each with a description and an example code snippet. Participants were then asked two questions per subcategory, using a 5-point Likert scale: (i) how frequently they encountered the inefficiency reported and (ii) how important they believed it was to address it. Additionally, the survey included an optional question inviting participants to report any inefficiency categories or subcategories not mentioned in the taxonomy based on their experience The survey remained open for $4$ weeks, and we sent a reminder email two weeks after the first call. Once the results for each inefficiency pattern were collected, we processed them to evaluate their frequency and relevance. This was achieved by aggregating the results using weighted averages of Likert scale values, following methodologies from similar studies [13]. Additionally, we complemented the analysis with comments from responders to validate the taxonomy.
IV Results
In this section, we present the results of the two RQs investigated in this study. First, we present our taxonomy of LLM-generated code inefficiencies, and then, we discuss the perceptions of our proposed taxonomy through our survey.
IV-A RQ1: Patterns of Inefficiencies in LLM-generated Code
IV-A1 Taxonomy
We organized our observations as a taxonomy of five categories with 19 subcategories. In the following, we explain each of the five categories with their respective subcategories. Examples of each can be found in our replication package (code snippets).
A. General Logic. This category encompasses inefficiencies related to the overall logic of a given task, including issues with understanding requirements and translating them into code.
  • Wrong Logic. It occurs when the generated code lacks meaningful logic, containing only comments, hard-coded values, returning method inputs, or using a simple pass statement. It also includes case-by-case implementations that fail to follow coherent logic. Additionally, it encompasses situations where all requirements are implemented incorrectly, with none being addressed properly.
  • Partially Wrong Logic Different from Wrong Logic, this subcategory occurs when some requirements are implemented correctly while others are either incorrectly implemented or not implemented at all.
  • Wrong Method Input. It occurs when the LLM generates a statement that is generally coherent with the logic but uses incorrect or inappropriate inputs. For example, passing a single element to the sorted method instead of a list may lead to unintended behavior.
  • Exception/Corner Case Handling. This occurs when the generated code lacks exception handling or fails to address corner or edge cases. Initially, this sub-category was merged with Partially Wrong. However, following survey participants’ suggestions, it was detached as a standalone subcategory.
B. Performance. This category groups inefficiencies that negatively impact the generated code’s performance.
  • A More Memory-Optimal Solution Exists (Memory). This subcategory captures cases where the same logic can be implemented in a more memory-efficient manner. For example, recursion may be avoided when intermediate results are stored on the call stack, potentially leading to memory overflow for deep recursion. Similarly, list comprehensions, which store all results in memory, can be replaced with generator expressions to enable lazy evaluation and reduce memory usage.
  • A more optimal solution exists (Time). It identifies the cases where an equivalent, more time-efficient algorithm exists compared to the LLM-generated code. In this subcategory, we report four additional subcategories:
    • Sub-Optimal Solution (Time Complexity): It occurs when the implemented logic can be replaced with another algorithm with lower time complexity. For instance, using Quick Sort, with a time complexity of $O(n\log n)$ on average, is generally more efficient than Insertion Sort, which has a time complexity of $O(n^{2})$ in the average and worst cases.
    • Unnecessary Steps. It refers to instructions in the code that do not affect or contribute to the final output. Eliminating such steps enhances the code’s overall efficiency without altering its functionality, such as unnecessary type conversion.
    • Redundant Steps. This occurs when an instruction essential to the logic is unnecessarily repeated. To avoid such cases, the given instruction could be executed once, and its output could be stored in a local variable instead of re-executing the same instruction.
    • Inefficient Iterative Block. This category focuses on inefficiencies within repetitive structures, such as loops and iteration management.
      + Iterative Block (Initialization Condition). It occurs when the initialization condition of a loop is overly broad and can be narrowed without affecting the method’s correctness. For example, when determining whether a number is prime, we can stop checking divisibility at $\sqrt{n}$ rather than iterating all the way up to $n$.
      + Iterative Block (Stopping Condition). It occurs when the stopping condition for a repetitive block is too broad and can be optimized without changing the intended functionality or when it lacks an early stopping condition when applicable. For instance, when searching for an element in a sorted list, instead of iterating through the entire list, we can stop early if the current element exceeds the target value.
C. Readability This category encompasses issues on the overall readability, clarity, and ease of comprehensibility of LLM-generated code snippets.
  • Confusing Variable Naming. It refers to variable names that can confuse, even if they do not result in runtime errors. It includes issues such as redefining existing names (e.g., using dict as a variable or method name) or variable shadowing, where the same variable name is used in different scopes.
  • A more readable method exists. This refers to cases where the code is difficult to understand or does not adhere to conventions. For example, the combination of list comprehension, map, and lambda functions, along with a for loop and multiple method calls, such as the use of this statement: return reduce(lambda x, y: x * y, [int(i) for i in str(n) if int(i) % 2 == 1]).
D. Maintainability. This category highlights inefficiencies in the LLM-generated codes that may negatively impact the maintainability of the code, possibly being intended to be integrated into larger systems.
  • Code Duplication. This occurs when identical or similar blocks of code appear in multiple places without adding new value to the main logic of the code, essentially resembling code cloning [41].
  • Comment Duplication. It occurs when repetitive comments are used excessively, identical or similar, without providing additional insight or explanations about the logic.
  • Conditional Block. It groups inefficiencies related to conditional blocks, i.e., if blocks. Although it does not impact the functionality, it is important to address it for better code clarity and easier testing/debugging [42]. While in our studied sample we did not observe foreach blocks, this subcategory may be extended to cover them as well.
    • Unnecessary Else. It takes place when an else statement is used unnecessarily within an if block and can be removed without affecting the program’s behavior. It usually happens when the preceding if or elif statements already contain a return or break, making the else unnecessary.
    • Unnecessary Conditional Block. It refers to cases where using a conditional block adds unnecessary complexity and can be simplified by directly using the output of a given condition in a return statement.
E. Errors Inefficiencies falling under Erros are related to the fundamental correctness of the code regardless of the logic being implemented. These issues can lead to runtime errors, unexpected behavior, or failure to execute the code properly.
  • Missing Import. It occurs when methods or constants from external modules/libraries are used without being properly imported. Missing imports can lead to errors or failures when the code attempts to use the functionality provided by those modules.
  • Missing Variable Declaration. It occurs when a variable is used without being properly declared, initialized, or passed as a parameter for a method, which may result in runtime errors.
  • Syntax Error. This occurs when the code contains a syntax error preventing it from being executed, such as improper use of keywords or incorrect indentations in Python.
IV-A2 Code Inefficiencies Patterns
TABLE I: (1) Frequency of inefficiencies in LLM-generated code, and (2) Survey-based taxonomy validation Results
(1)Frequency in of ineffeciencies in LLM-generated code(%)
(2) Survey-based Validation
(Weighted average)
Codellama Deeepseek-Coder CodeGemma Overall Popularity Relevance
1- General Logic 60.37 70.73 74.39 68.5 4.03
Wrong Logic 9.76 39.63 57.93 35.77 2.22 3.64
Partially Wrong Logic 47.56 27.44 15.24 30.08 3.17 3.88
Wrong Method Input 4.27 3.66 1.22 3.05 2.76 3.53
2- Performance 48.78 30.49 23.17 34.15 3.52
Sub-Optimal Solution (Memory) 26.83 17.07 12.2 18.7 3.07 3.12
Iterative Block (Initialization Condition) 0.61 0.61 0.61 0.61 2.69 3.19
Iterative Block (Stopping Condition) 3.66 1.22 2.44 2.44
Redundant Steps 9.76 3.05 3.05 5.28 2.64 3.05
Unnecessary Steps 4.27 2.44 0.61 2.44 3.03 3.03
Sub-optimal Time Complexity 23.78 17.68 14.02 18.5 2.93 3.12
3-Readability 6.71 6.1 1.22 4.67 3.22
Confusing Variable Naming 1.22 1.22 0.61 1.02 2.17 3
Sub-Readable Code Exists 5.49 4.88 0.61 3.66 2.79 3.05
4-Maintainability 29.88 12.8 20.73 21.14 3.31
Unnecessary Else 21.95 10.98 19.51 17.48 2.29 2.59
Unnecessary Conditional Block 3.66 1.83 3.66 3.05 2.62 2.78
Code Duplication 4.88 1.83 0.61 2.44 2.46 3.02
Comment Duplication 3.05 0 0.61 1.22 1.86 2.10
5-Errors 7.32 6.1 3.66 5.69 3.93
Syntax Error 0 2.44 2.44 1.63 2.29 3.79
Missing Variable Declaration 0.61 0 0.61 0.41 2.17 3.60
Missing Module Import 6.1 3.66 1.22 3.66 2.86 3.26
As previously discussed, our taxonomy of inefficiencies was derived from observed patterns in our studied sample using multi-labeling at both the category and subcategory levels. Regarding category frequency, most samples were labeled with a single category (56.5%), followed by those with two (24.59%) and none (9.96%). A similar trend was observed at the subcategory level, where the majority had one subcategory (54.07%), followed by two (18.29%) and three (11.38%). Notably, 33.54% of the samples exhibited characteristics spanning multiple categories, underscoring the complexity and overlap of inefficiency patterns across different code quality aspects. Table I presents the frequency of inefficiency categories and subcategories (2nd-5th columns). [^7] Since multi-labeling is allowed, the total may exceed 100% General Logic emerges as the most frequent category across all models (68.5% of samples). Among the selected models, CodeGemma, the least efficient, based on Top@k (leaderboard [33]), exhibited the highest rate of logic-related issues (74.39%), often producing code with no meaningful logic (57.93% of its cases). While CodeLlama, the most efficient, based on Top@K, showed fewer critical logic errors (60.37%), though it frequently generated partially correct solutions (47.56%), suggesting even advanced models may misinterpret requirements [43].
Performance inefficiencies ranked second (34.15%) with CodeLlama leading (48.78%). Sub-Optimal (Memory), Sub-Optimal (Time Complexity) and Redundant Steps were prevalent highlighting that even the most efficient model are prone to performance issues such as GPT-4 [44].
Regarding Maintainability issues varied significantly across models. DeepSeek-Coder, trained exclusively on code, exhibited the lowest maintainability flaws (12.8%), whereas CodeLlama and CodeGemma (trained on mixed code/natural language data) showed higher rates (29.88% and 20.73%, respectively). This discrepancy suggests that code-specific training may improve structural quality, while hybrid training may dilute maintainability focus. The prevalence of Unnecessary Else blocks (21.95% in CodeLlama) underscores a broader challenge: LLMs often generate verbose control flows, complicating long-term maintenance [45].
Readability (4.67%) and Errors (5.69%) were less frequent but non-trivial. Syntax errors were rare ($<2.44\%$), confirming LLMs’ strong syntactic grasp. However, Missing Imports (6.1% in CodeLlama) reveal persistent gaps in dependency resolution—a known limitation even in most-efficient models, GPT-4, [46].
Figure 2: Normalized Category Co-occurrence
IV-A3 Analysis of Co-occurrence Categories
To gain deeper insights into inefficiency relationships, we analyze how different categories co-occur across models. Figure 2 presents the normalized category co-occurrence heatmap across all models. The co-occurrence values are normalized row-wise, with each value divided by the frequency of the corresponding category. We observe strong inter-dependencies between inefficiencies in LLM-generated code. General Logic inefficiencies exhibit the highest co-occurrence rates with other categories, particularly Readability (0.74), Errors (0.64), and Maintainability (0.58). This suggests that flawed logic not only impacts correctness but also contributes to inefficiencies in code clarity, structure, and long-term maintainability. Similarly, Performance inefficiencies frequently co-occur with Readability (0.48) and Maintainability (0.46), highlighting that performance issues are often tied to code complexity and structure, rather than isolated algorithmic inefficiencies. Interestingly, Errors demonstrate weaker co-occurrences with other inefficiencies, with their strongest link being to Maintainability (0.36) indicating that while syntax and dependency errors are less common, they may still hinder maintainability efforts by requiring additional debugging and refactoring. Meanwhile, Readability and Maintainability are closely linked (0.35), reinforcing the idea that clearer code structures contribute to better long-term maintainability. These findings suggest that improving LLM-generated logic and performance could lead to broader improvements across multiple inefficiency categories, reducing the need for manual corrections and enhancing code usability in real-world development.
IV-B RQ2: Relevance of Identified Inefficiencies for Practitioners and Researchers
We kept the survey open for four weeks reaching $58$ responses, resulting in a response rate of $7.39$%, which aligns with typical rates observed in software engineering surveys [26, 47, 48, 49]. Here, we first present the demographic information about the participants and their experience with coding using LLMs. Second, we provide a detailed analysis of the reported frequency and perceived relevance of inefficiencies in LLM-generated code.
IV-B1 Demographic Information and Experience with Coding using LLMs
In our survey, all participants answered the demographic questions, obtaining the following answers.  [^8] Some percentages may exceed 100% as participants were able to select multiple options for some questions Our sample consisted of participants from both academia (60%) and industry (40%). Among academic participants, we had graduate, undergraduate students, and postdoctoral researchers/lecturers (29, 3, and 3, respectively). From the industry, we had 11 software developers, 7 ML/data science developers, 4 senior executive roles, and 1 Cybersecurity developer. About the experience, 37 participants had more than 5 years of experience, while 11 had between 3-5 years, and 10 had between 1-3 years. Regarding adopting LLMs to assist in coding, 98% of the participants report using them. Both proprietary and open-source LLMs are in use, with 68% relying exclusively on proprietary models. Among proprietary models, ChatGPT and Copilot are the most commonly used (89.66% and 29.31%, respectively). On the other hand, for open-source models, the most popular are Mistral, the Phi model family, and Qwen-Coder (5.17%, 3.45%, and 3.45%, respectively). Finally, participants also use a variety of programming languages, with the top five being Python, C/C++, JavaScript, Java, and Go (91.38%, 18.97%, 18.97%, 8.62%, and 3.45% respectively).
Before getting into the questions addressing our taxonomy, we asked participants about their general experience and faced issues in LLM-generated code. They reported inefficiencies that can be categorized into two main groups: user experience (outside the scope of our taxonomy) and code-related, which are directly relevant to our research. We believe such a diverse distribution across academia and industry, combined with a strong representation of experienced professionals, provides a well-balanced perspective on inefficiencies in LLM-generated code. The diversity in roles and expertise levels enhances the reliability of our findings by capturing insights from both research-oriented and real-world coding environments. We discuss the reported inefficiencies as follows.
User Experience-Related Inefficiencies
These inefficiencies focus on the challenges/hardness users face when interacting with LLMs, rather than the generated code. Overall, 6 participants reported that LLMs often perform poorly when tasked with debugging or modifying specific parts of the code, as also reported in previous studies [50]. In the same way, programming language and library familiarity also play a significant role in LLM efficiency. 5 Participants observed that LLMs tend to perform better with popular languages and widely used libraries, such as Python and NumPy, respectively, compared to less popular ones, like DHL and Tinkter [51]. Moreover, the programming language paradigm also matters; one participant noted that LLMs perform better with interpreted languages than with compiled ones. Participants also shared different experiences when using Chat-based LLMs, such as ChatGPT, compared to code assistants like Copilot. One participant mentioned preferring ChatGPT, reporting that Copilot made “stupid suggestions and distracted me”. Conversely, another participant found Copilot more useful for providing inline suggestions while maintaining overall visibility of the codebase findings [52].
Code-Related Inefficiencies
Participants reported various inefficiencies in the generated code that align with the focus of our taxonomy. 19 participants stated that LLMs often struggle to produce correct logic on the first attempt, even for simple tasks accompanied by detailed descriptions, requiring multiple iterations to achieve the desired outcome. P22 specifically noted LLMs may generate “unnecessarily long or complex (for simple tasks)” However, for complex tasks, as reported by P15, such as building a Flask application, LLMs tend to“oversimplification to the point of uselessness”, offering high-level steps or generalized code that lack specificity, performance efficiency, and functionality. Furthermore, 3 participants reported that handling exceptions and corner cases is often overlooked. All these reported issues fall under General Logic category of our taxonomy and were introduced spontaneously before participants were introduced to our taxonomy.
Another notable issue is the misalignment between LLM advancements and technological progress. 13 participants observed that LLMs often fail to account for updates in library versions, resulting in version tracking challenges. Examples include mixing methods from different library versions, using deprecated libraries, calling non-existent methods, or generating code incompatible with the specified library or API versions[46]. Some of these issues align with Wrong Method Input (General Logic) and Missing Import (Errors) sub-categories of our taxonomy. Participants—using various programming languages, like Python, C/C++, Java, JavaScript, Go, and Bash—also reported readability-related inefficiencies. They included poor code formatting, failure to adhere to user-defined coding styles or naming conventions, inconsistent or missing variable names, unused variables, and reliance on outdated coding practices. Such observations highlight that these issues may persist across different programming languages [16]. Additionally, 6 participants highlighted performance concerns, while others noticed maintainability issues, noting that the generated code is not always performant and often lacks robust, modular, or object-oriented architecture. Insufficient documentation and comments further exacerbate the maintainability challenges. These observations are reflected in the proposed Readability, Performance, and Maintainability categories of our taxonomy and aligned with findings in the literature [16, 14, 16]. 7 participants noted that context limitations were a significant factor in the issues they encountered. Their experiences with memorization in chat-based LLMs varied. Two participants found that LLMs had a limited context window and frequently forgot past interactions [53]. In contrast, another participant reported that LLMs tended to retain outdated context, stating that they needed to “start a new chat to get rid of hallucinations.”
IV-B2 Popularity and Relevance of Code Inefficiencies: Survey Perception
Table I (6th and 7th columns) summarizes the aggregated results of popularity and relevance of sub-categories of code inefficiency patterns on a scale of 1 to 5. The popularity score reflects how frequently participants faced each type of inefficiency, while relevance aims to indicate whether the participants consider such an inefficiency should be prioritized. Our findings reveal strong agreement between frequency of inefficiencies in the studied sample and popularity scores reported by participants.
General Logic inefficiencies were rated as the most relevant category (4.03/5). In particular Partially Wrong Logic stands out as the most relevant and popular subcategory encountered by the participants (3.88/5 and 3.17/5, respectively) and highly frequent in our studied sample (30.08%). Such an observation suggests that while LLMs generate code that often aligns with task descriptions, they frequently miss certain requirements, requiring manual iterations to refine the logic. Participants confirmed this emphasizing the need to repeatedly adjust LLM-generated code to achieve correctness. A challenge also noted in prior studies particularly in studies analyzing Copilot-generated code [10]. Although Wrong Logic is reported as the most frequent subcategory in our studied sample (35.77%), it receives a relatively low popularity score (2.22/5), which can be explained by the fact that participants mostly use more efficient propitiatory models with 98% relying on at least one such model.
Performance inefficiencies, occurring in 34.15% in our studied sample, were considered also highly relevant (3.52/5) and highly popular . The most concerning issues were Sub-Optimal Memory Usage and Sub-Optimal (Time Complexity) ( 3.0 for popularity, 3.12 for relevance), including Inefficient Iterative Blocks, which directly impact execution time. While participants generally do not rely on LLMs to produce optimized code, they still prefer efficiency and often manually refactor performance issues. Such a finding suggests that while correctness is prioritized, performance issues in LLM-generated code can still add to developers’ workload.
For Maintainability, participants reported a moderate popularity (2.3/5) but a high relevance score (3.31/5), indicating that while they encounter maintainability issues less frequently, they still consider them important. Unnecessary Else Statement, the most frequently observed maintainability issue in our studied sample (17.48%), had the lowest relevance score (2.59/5), as many developers considered it a minor concern that does not impact correctness. In contrast, Code Duplication, despite being the least frequent inefficiency (2.44%) and less popular (2.46/5), was rated more relevant (3.02/5) due to its impact on maintenance effort and refactoring complexity. This suggests that developers prioritize maintainability issues that directly increase long-term maintenance costs, even if they occur less often.
Readability inefficiencies were relatively rare (4.67% frequency) but still perceived as moderately relevant (3.22/5.0) despite low popularity scores. Such an observation highlights that even infrequent readability issues are critical to address, as developers typically resolve them manually rather than relying fully on automated fixes [54]. Finally, while Errors were the least frequent (5.69%) and the least popular ( 2.25/5) inefficiencies encountered by participants,they had a high relevance score (3.93/5.0). Participants emphasized that although rare, these errors must be addressed immediately, as they prevent the code from being compiled or executed.
The strong agreement between our studied sample and participant feedback validates the relevance of our results, with only minor differences observed. This alignment suggests that the inefficiencies identified in our analysis accurately reflect the challenges practitioners and researchers encounter when working with LLM-generated code.
IV-B3 LLM-generated Code Inefficiencies Taxonomy: Feedback and Completeness
For each sub-category, we gathered participants’ feedback (open questions). Regarding General Logic, participants indicated that Wrong Logic often occurs when tasks are complex or involve a new tech stack. For Partially Wrong Logic, participants acknowledged that they typically correct the code manually but emphasized that LLMs sometimes “generate buggy solutions, and debugging takes longer than writing the code from scratch”(P41). Additionally, some participants highlighted that partially correct logic can be difficult to detect, potentially leading to errors that remain unnoticed until runtime, representing a risk as reported by P2: “A junior coder who uses this could cause issues if pushed to production code without proper review”.
For Performance, participants from both academia and industry indicated that while they do not rely on LLMs for performant solutions, they prioritize bug-free code over-optimized but non-functional code, choosing to manually refine performance only when necessary.
Regarding Readability, participants expressed indifference to code clarity and comprehensibility as long as the code works and they can rely on their knowledge to understand it and improve it.
For Maintainability, participants noted that issues like code and comment duplication are less noticeable when using tools such as Copilot and Cursor AI. Concerning Unnecessary Else Statements, they felt the impact was minimal and were generally uninterested in addressing such inefficiencies. Participants also pointed out that Errors, though potentially causing runtime issues, are usually easy to detect using linters or IDEs and can be corrected without significant effort.
Participants acknowledge that while prompting can help address issues related to General Logic, Performance, and Readability, certain inefficiencies—such as Wrong Method Input—still require developer expertise and experience to be effectively resolved. Overall, participants agreed that the proposed taxonomy effectively captures inefficiencies in LLM-generated code, as evidenced by its high popularity and relevance scores. Many of these inefficiencies were already mentioned by participants before they were introduced to the taxonomy (IV-B1), demonstrating strong coverage. However three participants suggested introducing a new sub-category for Exception and Corner Case Handling, as these issues were not explicitly covered.
V Related Work
While prior research has analyzed various quality aspects and revealed inefficiencies in LLM-generated code, to the best of our knowledge, no study has systematically examined, categorized, and constructed a comprehensive taxonomy of inefficiencies based on observed patterns in generated code. Our work addresses this gap by defining and classifying inefficiencies in LLM-generated code, providing new insights into their nature and prevalence.
Zheng et al. [16] studied LLM-generated code quality across multiple aspects, including readability, maintainability, correctness, and efficiency. Their findings suggest that LLMs exhibit a distinct coding style that does not always align with standard best practices, leading to concerns about long-term maintainability. Similarly, Licorish et al. [17] found that human-written code adheres more closely to established coding conventions than GPT-4-generated code. Dou et al. [14] further observed that while LLMs tend to generate shorter solutions for complex problems, these solutions often exhibit higher structural complexity, making them more difficult to maintain and debug.
Maintaining and optimizing LLM-generated code remains another critical challenge. Ouedraogo et al. [55] identified multiple test smells, such as Useless Tests and Assertion Roulette, in unit tests generated by LLMs, highlighting a lack of test quality. Similarly, Siddiq et al. [11] reported that GPT-generated code often suffers from improper documentation, unused or undefined variables, and security vulnerabilities, which can increase maintenance efforts. Our findings complement these studies by showing that such inefficiencies extend beyond isolated instances and are widespread across LLM-generated code.
Security risks in LLM-generated code have also been extensively studied. Fu et al. [39] examined vulnerabilities in Copilot-generated code and identified 43 Common Weakness Enumeration (CWE) issues, including eight from the CWE Top 25, reinforcing concerns that LLM-generated code may introduce significant security flaws if not carefully reviewed.
Researchers have also investigated the occurrence of bugs in LLM-generated outputs. Moradi Dakhel et al. [10] evaluated Copilot’s performance on fundamental algorithmic problems and found that while it can generate functional solutions, some are buggy solutions. Tambon et al. [13] analyzed buggy open-source LLM-generated code and introduced a taxonomy of bug patterns, while Dou et al. [14] categorized the root causes of common bugs, emphasizing issues related to model limitations and training data.
We observe that our proposed taxonomy intersects with prior taxonomies primarily in two categories, along with their respective subcategories: General Logic and Errors, both of which contribute to buggy code. However, inefficiencies related to Performance, Readability, and Maintainability are largely overlooked in previous studies. This is because code exhibiting these inefficiencies is not necessarily considered buggy but may still impact software quality and performance. Our taxonomy addresses this gap by providing a more comprehensive classification that captures inefficiencies beyond mere correctness, offering new insights into the challenges of integrating LLM-generated code into real-world development.
VI Threats to Validity
Construct Validity The process of collecting and labeling generated code may influence our study’s results and findings. While our methodology aligns with existing works proposing taxonomies in SE [13, 26], we mitigate such a risk by clearly explaining our approach to ensure external validation. In the absence of LLM-generated code inefficiency categorization, we recognized the risk of introducing bias into our classification. To address this, we adopted an open coding procedure and utilized different LLM models to enhance the robustness and reliability of our findings. Additionally we surveyed practitioners and researchers to validate the completeness of our taxonomy.
Internal Validity A key internal threat to validity is the potential bias in the manual labeling of LLM-generated code. To mitigate such a threat, the first two authors—a senior Master’s student and a postdoctoral fellow with expertise in SE research and Python programming—independently labeled the generated code snippets. The inter-rater agreement at the category level is 0.846, which aligns with similar studies in software engineering [26, 13]. Additionally, to further validate our findings, we surveyed 58 participants from both academia and industry, ensuring a broader perspective on the identified inefficiencies. Furthermore, we leveraged an LLM as a judge to assess generated code snippets, providing an additional layer of validation and reducing potential human bias in the labeling process.
External Validity One external threat to validity concerns the selection of LLMs used to generate code. In this study, we utilize three open-source LLMs: CodeLlama, DeepSeek-Coder, and CodeGemma, all of which have demonstrated efficiency and have been employed in prior studies to leverage the capabilities of LLMs for code generation [56, 57, 58] Another potential threat is regarding the complexity of the benchmark tasks, as they may not fully represent real-world programming scenarios, which could limit the generalizability of the inefficiencies identified. To mitigate this threat, we use the HumanEval+ dataset, a widely adopted benchmark for evaluating LLM-generated code [28, 29, 31]. Such a benchmark consists of human-crafted algorithmic tasks, which, while not encompassing the full breadth of software development tasks, represent a fundamental component of real-world programming. Furthermore, our study focuses exclusively on Python code, the most widely used programming language. However, this scope introduces a limitation, as we may have overlooked inefficiencies that are specific to other programming languages. Future work could extend our analysis to additional programming languages to enhance generalizability. Another critical threat is the long-term relevance of our proposed taxonomy, given the rapid evolution of LLMs. As these models continue to improve, some inefficiencies may become less prevalent, while new ones may emerge. To address this threat, future research should periodically reassess and refine the taxonomy to ensure its continuous applicability in evaluating LLM-generated code.
Conclusion Validity Conclusions drawn in this study may have limitations regarding taxonomy incompleteness. Aiming to address such a threat, we conducted a manual inspection of 492 generated codes by three different LLM models. To reduce the potential error, each sample is labeled separately by the two labelers and subsequently discussed. Finally, the results are validated with a survey. We provided a replication package to facilitate the reproducibility of our findings.
VII Conclusion
In this study, we systematically investigated inefficiencies in LLM-generated code, introduced and validated through a survey a taxonomy categorizing these issues into 5 categories: General Logic, Performance, Readability, Maintainability, and Errors, with 19 specific subcategories. Our results showed that logic and performance inefficiencies are the most frequent and relevant often co-occurring and influencing software maintainability. Our findings highlight the need for LLMs to improve reasoning and optimization capabilities to mitigate other inefficiencies. The taxonomy provides a structured basis for evaluating LLM-generated code, facilitating efficiency comparisons, and guiding future improvements in LLMs for code generation.